OpenStudy (anonymous):

General solution to differential equation dy/dx = xy^2?

random231 (random231):

where are you stuck?

OpenStudy (anonymous):

This is what i have so far... is this right? \[dy \frac{ 1 }{ y ^{2} } = x dx\] \[\int\limits y ^{-2} dy = \int\limits x dx\] \[\frac{ y ^{-1} }{ -1 } = \frac{ x ^{2} }{ 2 } + C\] \[\frac{ 1 }{ y } = -\frac{ x ^{2} }{ x ^{2} } - C\] \[y = \frac{ 2 }{ -x ^{2} } + C\]

OpenStudy (anonymous):

Oh wait, there's a mistake in my equation

random231 (random231):

yep x^2/x^2

OpenStudy (anonymous):

The second to last equation should be -x^2 / 2 - C

OpenStudy (anonymous):

what about everything else, is my final answer wrong?

random231 (random231):

not exactly correct!

OpenStudy (anonymous):

hmm.. please give me a clue!

random231 (random231):

when you are inverting the 1/y to y,,, on the right side you cannot invert x^2/2 + c to 2/x^2 +C

random231 (random231):